# CHEMOUT: CHEMical complexity in star-forming regions of the OUTer Galaxy

### v1

@@ 4.1 Detection summary
A comprehensive molecular identification has been performed over the whole sample to identify the species detected in the greatest number of targets. This allows us to analyze abundance gradients as a function of Galactocentric distance with the highest possible statistics. Over 30 molecular species have been detected, and the most frequently detected ones are HCN (39/39), HCO+ (39/39), c-C3H2 (38/39), H13CO+ (30/39), HCO (29/39), and SO (25/39). The detection statistics over these species have already been shown in Fontani et al. (2022). We also included H2CO (18/18) and CH3OH (17/18) among the species considered. The analysis of the HCO, H2CO, and CH3OH gradients was conducted in Fontani et al. (2022b); however, here, we present updated trends based both on the revised fits performed with MADCUBA (Sect. 4.1.1) and on our new H2 column density estimates from NIKA2 observations. The spectroscopic parameters of the transitions detected in these molecules are displayed in Table 3. In Appendix C, the best-fit parameters obtained from the analysis with MADCUBA, and the spectra of the fitted transitions for HCN, H13CO+, and SO are shown in detail. The spectra of c-C3H2 and HCO+ are shown in Fontani et al. (2022), and the ones of HCO, H2CO, and CH3OH in Fontani et al. (2022b). The results regarding the molecular column densities (e.g., their Galactocentric gradients) are provided in Appendix D.
@@
In Table 5, the $N_{\text{H}_{2}}$ values calculated assuming a gas-to-dust ratio from Giannetti et al. (2017) are presented. These values are used hereafter for the analysis. All the data, results, and discussions concerning fractional abundances with respect to H2 are based on these estimates, rather than on those obtained assuming a fixed $\gamma=100$. The NIKA2 dust continuum maps and the Galactocentric gradients for H2 column densities are shown in Appendix B.
@@
Six sources (e.g., 19383+2711, 19571+3113, WB89-006, WB89-060, WB89-080, and WB89-380) show multiple velocity features as reported also in Fontani et al. (2022). Four of them (19383+2711, 19571+3113, WB89-006, and WB89-380) exhibit double-peak emission in most of the observed lines. Because the peak velocities of these two features are consistent in the different transitions in which they are detected, they probably arise from two clumps within the telescope beam that have an angular separation smaller than the beam itself. This is confirmed by the NIKA2 maps (Figs. 4 - 7), for 19383+2711 and 19571+3113, showing two different emission components in the beam of the IRAM 30m telescope. WB89-006 and WB89-380 do not show a clear double peak in the continuum maps, but the two clumps could be non-resolved with NIKA2 or aligned along the line of sight. In principle, each of these clumps may have distinct physical and chemical properties and will therefore be analyzed separately in the following sections. For these, the peak velocity were estimated using the $V_{\text{LSR}}$ of the c-C3H2 $J_{K_{a},K_{c}}=2_{1,2}-1_{0,1}$ line, and we assumed the same Galactic longitude $\ell$, for both clumps, to estimate their Galactocentric radius $R_{\text{GC}}$ and heliocentric distance $d$. The remaining two sources (WB89-060 and WB89-080) are characterized by a double peak only in some more intense and potentially optically thick lines (e.g., HCN, HCO+) and do not show coherence in the peak velocities. Consequently, in these cases, the double-peaked line shapes are more likely due to high-optical depth effects, like (self-)absorption, rather than to multiple velocity features. These have thus been analyzed as single sources, and the absorbed lines discarded from the analysis. Moreover, some transitions showed non-Gaussian, high-velocity wings. We decided to limit our analysis to the narrower component of the molecular emissions related to the line emission around the intensity peak, and thus arising from the bulk emission of the cores.
@@
Table 3: Rest frequencies and other spectroscopic parameters of the studied and detected molecular transitions
@@ 4.1.1 Formaldehyde and methanol updated fits
The fits for H2CO and CH3OH have been revised due to an incorrect assumption in Fontani et al. (2022b), which originated from a conversion between $T_{\mathrm{A}}^{*}$ and $T_{\rm MB}$. Since the beam efficiency $\eta_{\text{eff}}$ of the telescope was not explicitly specified in Bernal et al. (2021), the values were taken from Appendix C.3 of the ARO 12m user manual; however, outdated efficiencies were inadvertently used. Updated efficiencies333from https://aro.as.arizona.edu/?q=beam-efficiencies have now been adopted to estimate the $T_{\mathrm{MB}}$ of these two species. For H2CO at 2 mm we divided $T_{\mathrm{A}}^{*}$ by 0.81, and for CH3OH at 3 mm we divided by 0.91. The source sizes for H2CO and CH3OH were re-estimated using Eq. 3 in Fontani et al. (2002). Consequently, the MADCUBA fits were recalculated, and the resulting parameters are reported in Appendix C. The excitation temperatures of CH3OH are listed in Table 5.
@@ 4.2 Fractional abundances
Table 4: Fractional abundance, $X_{\text{mol}}$, w.r.t. H2 of HCN, HCO+, c-C3H2, H13CO+, HCO, SO, H2CO and CH3OH.444
@@
The plots showing $X_{\text{mol}}$ against $R_{\text{GC}}$ for each molecule are shown in Fig. 1. For comparison, the fractional abundances calculated using a gas-to-dust ratio $\gamma=100$ are shown too. A linear fit is performed for each plot, and the Pearson correlation coefficient [^5] This coefficient measures the strength and direction of the linear relationship between two variables, X and Y, and it is calculated as the covariance of X and Y divided by the product of their standard deviations. It ranges from 1-1 to +1+1, where +1+1 indicates perfect positive correlation, 0 indicates no linear correlation, and 1-1 indicates perfect negative correlation., $\rho$, is calculated to inspect possible (anti-) correlations. The linear fit applied to the data of fractional abundances, calculated using a gas-to-dust ratio from Giannetti et al. (2017), gives:
$[\text{HCN}]$ $=(-0.105\pm 0.013)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(8.79\pm 0.20)\ \mathrm{dex}$ (3)
$[\text{HCO}^{+}]$ $=(-0.078\pm 0.018)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(9.50\pm 0.28)\ \mathrm{dex}$ (4)
$[\textit{c}\text{-C}_{3}\text{H}_{2}]$ $=(-0.058\pm 0.021)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(9.95\pm 0.32)\ \mathrm{dex}$ (5)
$[\text{H}^{13}\text{CO}^{+}]$ $=(-0.039\pm 0.025)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(11.32\pm 0.38)\ \mathrm{dex}$ (6)
$[\text{HCO}]$ $=(-0.057\pm 0.026)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(9.83\pm 0.41)\ \mathrm{dex}$ (7)
$[\text{SO}]$ $=(-0.080\pm 0.023)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(8.98\pm 0.36)\ \mathrm{dex}$ (8)
$[\text{H}_{2}\text{CO}]$ $=(-0.103\pm 0.050)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(8.42\pm 0.75)\ \mathrm{dex}$ (9)
$[\text{CH}_{3}\text{OH}]$ $=(-0.146\pm 0.054)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(7.67\pm 0.82)\ \mathrm{dex}$ (10)
@@
Figure 1: Galactocentric gradients of fractional abundances w.r.t. H2, $X_{\text{mol}}$. The plot shows the trends for HCN (a), HCO+(b), c-C3H2(c), H13CO+(d), HCO (e), SO (f), H2CO (g), and CH3OH (h) as a function of the Galactocentric radius ($R_{\text{GC}}$). The light blue data represent the abundances calculated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue data illustrate those estimated using a non-constant gas-to-dust ratio ($\gamma=\gamma(R_{\text{GC}})$ from Giannetti et al. 2017). For both datasets, the linear regression results are shown as the light blue and dark blue lines, respectively. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. The upper limit values are represented with triangles. The gradients of the elemental abundances of carbon (C, and 13C), nitrogen (N), oxygen (O), and sulfur (S), as reported by Méndez-Delgado et al. (2022), are plotted as dashed lines. The product of the parent elements of the species is represented, in each subplot, by a black dash–dot line. All elemental trends are plotted with a vertical offset to align with the starting point of the molecular gradients derived using $\gamma(R_{\text{GC}})$, to facilitate comparison of their slopes with the linear fit of the molecular abundances. In the upper-right side of each subplot, the Pearson correlation coefficient, $\rho$, is shown (estimated only for the abundances estimated using the non-constant gas-to-dust ratio).
@@ 4.3 Line widths
From the fit of the lines, we also obtained the full line widths at half maximum (FWHM). The measured FHWMs for the molecules studied are in the ranges shown in Table 2. The mean values of line width estimated for each molecule have been used in the estimation of the upper limits of the column densities, as an educated guess ($2.1\pm 0.6\penalty 10000\$km s-1 for HCN; $1.8\pm 0.9\penalty 10000\$km s-1 for HCO+; $2.1\pm 1.0\penalty 10000\$km s-1 for c-C3H2; $2.1\pm 1.0\penalty 10000\$km s-1 for H13CO+; $2.2\pm 1.0\penalty 10000\$km s-1 for HCO; $2.0\pm 0.8\penalty 10000\$km s-1 for SO; $2.3\pm 0.7\penalty 10000\ \mathrm{km\ s^{-1}}$ for H2CO; $1.9\pm 0.8\penalty 10000\ \mathrm{km\ s^{-1}}$ for CH3OH).
@@
Figure 2: Galactocentric gradients of line widths, FWHM. The plot shows the trends for HCN (a), HCO+(b), c-C3H2(c), H13CO+(d), HCO (e), SO (f), H2CO (g), and CH3OH (h) as a function of the Galactocentric radius ($R_{\text{GC}}$). The linear fit results are shown as the dark blue lines. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. In the upper-right side of each subplot, the Pearson correlation coefficient, $\rho$, is shown.
@@ 5.1 Galactocentric gradients of fractional abundances
The trends obtained for the molecular abundance (w.r.t. H2) have been compared with the Galactocentric abundance gradients of carbon (both 12C and 13C), oxygen, nitrogen, and sulfur. These are constituting elements of the studied species, and the gradients have been determined in Méndez-Delgado et al. (2022) through the analysis of the spectra of Galactic H II regions. The Galactocentric gradient of 13C has been estimated by coupling the carbon gradient with the Galactocentric trend of the $^{13}\text{C}/^{12}\text{C}$ ratio from Yan et al. (2023). Between 8.34 kpc, the Galactocentric distance of the Sun, and 24 kpc, the distance of the most distant source in our sample, the abundance ratios [C/H], [13C/H], [N/H], [O/H], and [S/H] decrease by factors of approximately $\sim 16$, $\sim 36$, $\sim 10$, $\sim 5$, and $\sim 4$, respectively. Extrapolating the gradients derived in the OG (Eqs. 38) to the Solar circle, over the same range of $R_{\text{GC}}$, we find that the abundance of the studied species, i.e. HCN, HCO+, c-C3H2, H13CO+, HCO, SO, H2CO, and CH3OH decrease by a factor $\sim 44$, $\sim 17$, $\sim 8$, $\sim 4$, $\sim 8$, $\sim 18$, $\sim 41$, and $\sim 194$
@@
The decreasing rates of HCN, HCO+, c-C3H2, HCO, H2CO, and CH3OH are consistent, within the error bars, with that of [C/H]. This implies that, considering the dispersion and uncertainties in estimating these quantities, the fractional abundances of these carbon-bearing species scale with the elemental abundance of carbon ([C/H]) between 8.34 kpc and 24 kpc. This suggests that the efficiency in forming these molecules, which scales with the availability of the parent element C, is at least as high as in the local Galaxy.
@@
In contrast, the decreasing rate of SO appears steeper than that of [S/H]. For SO, the higher decreasing rate with respect to [S/H] indicates a lower efficiency of the formation of these molecules in the OG with respect to the local Galaxy. We speculate that a steeper decreasing trend for SO, which is the best shock tracer among those that we analyze here, could indicate that, in the OG, outflows have a reduced influence on the gas composition compared to the local Galaxy. This result would be consistent with the decreasing line width with $R_{\text{GC}}$ (Sect. 4.3). Another shock tracer is CH3OH, which, although its decreasing trend is marginally consistent with that of carbon, exhibits the most pronounced reduction among the species investigated in the OG. The steep decreasing trend of CH3OH could also arise from a less efficient grain-surface chemistry in the OG. The results concerning HCO, H2CO, and CH3OH, obtained with the updated abundances (and distances), confirm the ones of Fontani et al. (2022b). In Fontani et al. (2024), a correlation between the abundances of CH3OH and SO was found, which is confirmed over the whole CHEMOUT sample, although no correlation between the line widths of these two species is observed, hinting that they are not tracing the same gas components.
@@
The decreasing rate of H13CO+, although lower than that of [13C/H], aligns well with the decreasing rate observed for oxygen. This correlation is not observed for HCO+, whose decreasing rate is consistent with carbon rather than oxygen. This discrepancy could suggest a possible revision of the Galactocentric gradient of the 13C/12C ratio, with a shallower trend at high $R_{\text{GC}}$, in agreement with the Galactic chemical evolution discussed in Colzi et al. (2022). Moreover, the detailed analysis of the 13C/12C ratios from the HCO+ and H13CO+ lines will be the subject of a forthcoming paper (Colzi et al. in prep.).
Therefore, the results regarding c-C3H2, HCO, and especially H13CO+ lead to the puzzling conclusion of observing more carbon-bearing molecules than the available carbon atoms in the OG, being their decreasing rates lower than that of [C/H] (or [13C/H]).

### v2

@@ 4.1 Detection summary
A comprehensive molecular identification was performed over the whole sample to identify the species detected in the greatest number of targets. This allowed us to analyze abundance gradients as a function of galactocentric distance with the highest possible statistics. Over 30 molecular species have been detected, and the most frequently detected ones are HCN (39/39), HCO+ (39/39), c-C3H2 (38/39), H13CO+ (30/39), HCO (29/39), and SO (25/39). The detection statistics over these species are already listed in fontani2022chemout. We also included H2CO (18/18) and CH3OH (17/18) among the species considered. The analysis of the HCO, H2CO, and CH3OH gradients was conducted by fontani2022chemoutII; however, here, we present updated trends based both on the revised fits performed with MADCUBA (Sect. 4.1.1) and on our new H2 column-density estimates from NIKA2 observations. The spectroscopic parameters of the transitions detected in these molecules are displayed in Table 3. In Appendix C, the best-fit parameters obtained from the analysis with MADCUBA, and the spectra of the fit transitions for HCN, H13CO+, and SO are shown in detail. The spectra of c-C3H2 and HCO+ are shown in fontani2022chemout, and the ones of HCO, H2CO, and CH3OH in fontani2022chemoutII. The results regarding the molecular column densities (e.g., their galactocentric gradients) are provided in Appendix D.
@@
In Table 5, the $N_{\text{H}_{2}}$ values calculated assuming a gas-to-dust ratio from giannetti2017galactocentric are presented. These values are used hereafter for the analysis. All the data, results, and discussions concerning fractional abundances with respect to H2 are based on these estimates, rather than on those obtained assuming a fixed $\gamma=100$. The NIKA2 dust continuum maps and the galactocentric gradients for H2 column densities are shown in Appendix B.
@@
Six sources (19383+2711, 19571+3113, WB89-006, WB89-060, WB89-080, and WB89-380) show multiple velocity features also reported in fontani2022chemout. Four of them (19383+2711, 19571+3113, WB89-006, and WB89-380) exhibit double-peak emission in most of the observed lines. Because the peak velocities of these two features are consistent in the different transitions in which they are detected, they probably arise from two clumps within the telescope beam that have an angular separation smaller than the beam itself. This is confirmed by the NIKA2 maps (Figs. 4 - 6), for 19383+2711 and 19571+3113, showing two different emission components in the beam of the IRAM 30 m telescope. WB89-006 and WB89-380 do not show a clear double peak in the continuum maps, but the two clumps could be non-resolved with NIKA2 or aligned along the line of sight. In principle, each of these clumps may have distinct physical and chemical properties and will therefore be analyzed separately in the following sections. For these, the peak velocities were estimated using the $V_{\text{LSR}}$ of the c-C3H2 $J_{K_{a},K_{c}}=2_{1,2}-1_{0,1}$ line, and we assumed the same Galactic longitude, $\ell$, for both clumps, to estimate their galactocentric radius, $R_{\text{GC,}}$ and heliocentric distance, $d$. The remaining two sources (WB89-060 and WB89-080) are characterized by a double peak, but only in some more intense and potentially optically thick lines (e.g., HCN and HCO+) and do not show coherence in the peak velocities. Consequently, in these cases, the double-peaked line shapes are more likely due to high-optical-depth effects, such as (self-)absorption, rather than multiple velocity features. These have thus been analyzed as single sources, and the absorbed lines discarded from the analysis. Moreover, some transitions showed non-Gaussian, high-velocity wings. We decided to limit our analysis to the narrower component of the molecular emissions related to the line emission around the intensity peak, and thus arising from the bulk emission of the cores.
@@
Table 3: Rest frequencies and other spectroscopic parameters of the studied and detected molecular transitions.
@@ 4.1.1 Formaldehyde and methanol updated fits
The fits for H2CO and CH3OH have been revised due to an incorrect assumption in fontani2022chemoutII, which originated from a conversion between $T_{\mathrm{A}}^{*}$ and $T_{\rm MB}$. Since the beam efficiency $\eta_{\text{eff}}$ of the telescope was not explicitly specified in bernal2021methanol, the values were taken from Appendix C.3 of the ARO 12m user manual; however, outdated efficiencies were inadvertently used. Updated efficiencies [^3] From https://aro.as.arizona.edu/?q=beam-efficiencies. have now been adopted to estimate the $T_{\mathrm{MB}}$ of these two species. For H2CO at 2 mm we divided $T_{\mathrm{A}}^{*}$ by 0.81, and for CH3OH at 3 mm we divided by 0.91. The source sizes for H2CO and CH3OH were re-estimated using Eq. 3 in fontani2002structure. Consequently, the MADCUBA fits were recalculated, and the resulting parameters are reported in Appendix C. The excitation temperatures of CH3OH are listed in Table 5.
@@ 4.2 Fractional abundances
Table 4: Fractional abundance, $X_{\text{mol}}$, with respect to H2 of HCN, HCO+, c-C3H2, H13CO+, HCO, SO, H2CO, and CH3OH.444
@@
The plots showing $X_{\text{mol}}$ against $R_{\text{GC}}$ for each molecule are shown in Fig. 1. For comparison, the fractional abundances calculated using a gas-to-dust ratio $\gamma=100$ are also shown. A linear fit is performed for each plot, and the Pearson correlation coefficient [^5] This coefficient measures the strength and direction of the linear relationship between two variables, X and Y, and it is calculated as the covariance of X and Y divided by the product of their standard deviations. It ranges from 1-1 to +1+1, where +1+1 indicates perfect positive correlation, 0 indicates no linear correlation, and 1-1 indicates perfect negative correlation., $\rho$, is calculated to inspect possible (anti-)correlations. The linear fit applied to the data of fractional abundances, calculated using a gas-to-dust ratio from giannetti2017galactocentric, gives:
$[\text{HCN}]$ $=(-0.105\pm 0.013)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(8.79\pm 0.20)\ \mathrm{dex},$ (3)
$[\text{HCO}^{+}]$ $=(-0.078\pm 0.018)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(9.50\pm 0.28)\ \mathrm{dex},$ (4)
$[\textit{c}\text{-C}_{3}\text{H}_{2}]$ $=(-0.058\pm 0.021)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(9.95\pm 0.32)\ \mathrm{dex},$ (5)
$[\text{H}^{13}\text{CO}^{+}]$ $=(-0.039\pm 0.025)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(11.32\pm 0.38)\ \mathrm{dex},$ (6)
$[\text{HCO}]$ $=(-0.057\pm 0.026)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(9.83\pm 0.41)\ \mathrm{dex},$ (7)
$[\text{SO}]$ $=(-0.080\pm 0.023)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(8.98\pm 0.36)\ \mathrm{dex},$ (8)
$[\text{H}_{2}\text{CO}]$ $=(-0.103\pm 0.050)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(8.42\pm 0.75)\ \mathrm{dex},$ (9)
$[\text{CH}_{3}\text{OH}]$ $=(-0.146\pm 0.054)\ \mathrm{dex\ kpc}^{-1}\cdot R_{\text{GC}}-(7.67\pm 0.82)\ \mathrm{dex.}$ (10)
@@
Figure 1: Galactocentric gradients of fractional abundances with respect to H2, $X_{\text{mol}}$. The plot shows the trends for HCN (a), HCO+(b), c-C3H2(c), H13CO+(d), HCO (e), SO (f), H2CO (g), and CH3OH (h) as a function of the galactocentric radius ($R_{\text{GC}}$). The light blue data represent the abundances calculated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue data illustrate those estimated using a non-constant gas-to-dust ratio ($\gamma=\gamma(R_{\text{GC}})$ from giannetti2017galactocentric). For both datasets, the linear regression results are shown as the light blue and dark blue lines, respectively. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. The upper-limit values are represented with triangles. The gradients of the elemental abundances of carbon (C and 13C), nitrogen (N), oxygen (O), and sulfur (S), as reported by mendez2022gradients, are plotted as dashed lines. The product of the parent elements of the species is represented, in each subplot, by a dash–dotted black line. All elemental trends are plotted with a vertical offset to align with the starting point of the molecular gradients derived using $\gamma(R_{\text{GC}})$, to facilitate comparison of their slopes with the linear fit of the molecular abundances. In the upper-right side of each subplot, the Pearson correlation coefficient, $\rho$, is shown (estimated only for the abundances estimated using the non-constant gas-to-dust ratio).
@@ 4.3 Line widths
From the fit of the lines, we also obtained the full widths at half maximum (FWHMs). The measured FWHMs for the molecules studied are in the ranges shown in Table 2. The mean values of line width estimated for each molecule have been used in the estimation of the upper limits of the column densities, as an educated guess ($2.1\pm 0.6\penalty 10000\$km s-1 for HCN; $1.8\pm 0.9\penalty 10000\$km s-1 for HCO+; $2.1\pm 1.0\penalty 10000\$km s-1 for c-C3H2; $2.1\pm 1.0\penalty 10000\$km s-1 for H13CO+; $2.2\pm 1.0\penalty 10000\$km s-1 for HCO; $2.0\pm 0.8\penalty 10000\$km s-1 for SO; $2.3\pm 0.7\penalty 10000\ \mathrm{km\ s^{-1}}$ for H2CO; $1.9\pm 0.8\penalty 10000\ \mathrm{km\ s^{-1}}$ for CH3OH).
@@
Figure 2: Galactocentric gradients of line widths, FWHM. The plot shows the trends for HCN (a), HCO+(b), c-C3H2(c), H13CO+(d), HCO (e), SO (f), H2CO (g), and CH3OH (h) as a function of the galactocentric radius ($R_{\text{GC}}$). The linear fit results are shown as dark blue lines. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. In the upper right side of each subplot, the Pearson correlation coefficient, $\rho$, is shown.
@@ 5.1 Galactocentric gradients of fractional abundances
The trends obtained for the molecular abundance (with respect to H2) were compared with the galactocentric abundance gradients of carbon (both 12C and 13C), oxygen, nitrogen, and sulfur. These are constituting elements of the studied species, and the gradients were determined in mendez2022gradients through the analysis of the spectra of Galactic H II regions. The galactocentric gradient of 13C was estimated by coupling the carbon gradient with the galactocentric trend of the $^{13}\text{C}/^{12}\text{C}$ ratio from yan2023direct. Among 8.34 kpc, the galactocentric distance of the Sun, and 24 kpc, which is the distance of the most distant source in our sample, the abundance ratios [C/H], [13C/H], [N/H], [O/H], and [S/H] decrease by factors of approximately $\sim 16$, $\sim 36$, $\sim 10$, $\sim 5$, and $\sim 4$, respectively. Extrapolating the gradients derived in the OG (Eqs. 38) to the Solar circle, over the same range of $R_{\text{GC}}$, we find that the abundance of the studied species, i.e., HCN, HCO+, c-C3H2, H13CO+, HCO, SO, H2CO, and CH3OH decrease by a factor of $\sim 44$, $\sim 17$, $\sim 8$, $\sim 4$, $\sim 8$, $\sim 18$, $\sim 41$, and $\sim 194.$
@@
The decreasing rates of HCN, HCO+, c-C3H2, HCO, H2CO, and CH3OH are consistent, within the error bars, with that of [C/H]. This implies that, considering the dispersion and uncertainties in estimating these quantities, the fractional abundances of these carbon-bearing species scale with the elemental abundance of carbon ([C/H]) between 8.34 kpc and 24 kpc. This suggests that the efficiency in forming these molecules, which scales with the availability of the parent element, C, is at least as high as in the local Galaxy.
@@
In contrast, the decreasing rate of SO appears steeper than that of [S/H]. For SO, the higher decreasing rate with respect to [S/H] indicates a lower efficiency of the formation of these molecules in the OG with respect to the local Galaxy. We speculate that a steeper decreasing trend for SO, which is the best shock tracer among those that we analyzed in this work, could indicate that, in the OG, outflows have a reduced influence on the gas composition compared to the local Galaxy. This result would be consistent with the decreasing line width with $R_{\text{GC}}$ (Sect. 4.3). Another shock tracer is CH3OH, which, although its decreasing trend is marginally consistent with that of carbon, exhibits the most pronounced reduction among the species investigated in the OG. The steep decreasing trend of CH3OH could also arise from a less efficient grain-surface chemistry in the OG. The results concerning HCO, H2CO, and CH3OH, obtained with the updated abundances (and distances), confirm the ones of fontani2022chemoutII. In fontani2024chemout, a correlation between the abundances of CH3OH and SO was found, which is confirmed over the entire CHEMOUT sample, although no correlation between the line widths of these two species is observed, hinting that they are not tracing the same gas components.
@@
The decreasing rate of H13CO+, although lower than that of [13C/H], aligns well with the decreasing rate observed for oxygen. This correlation is not observed for HCO+, whose decreasing rate is consistent with carbon rather than oxygen. This discrepancy could suggest a possible revision of the galactocentric gradient of the 13C/12C ratio, with a shallower trend at high $R_{\text{GC}}$, in agreement with the Galactic chemical evolution discussed in colzi2022chemout. Moreover, the detailed analysis of the 13C/12C ratios from the HCO+ and H13CO+ lines will be the subject of a forthcoming paper (Colzi et al. in prep.). Therefore, the results regarding c-C3H2, HCO, and especially H13CO+ lead to the puzzling conclusion of observing more carbon-bearing molecules than the available carbon atoms in the OG, their decreasing rates being lower than that of [C/H] (or [13C/H]).